OpenStudy (anonymous):

solve: log(underscore:x)9 = 1/2

OpenStudy (lgbasallote):

_ <--underscore (shift + minus sign)

OpenStudy (lgbasallote):

anyway... \[\log_x 9 = \frac 12\]

OpenStudy (lgbasallote):

right?

OpenStudy (anonymous):

yes igbasallote

OpenStudy (anonymous):

that's the equation

OpenStudy (lgbasallote):

\[\log_x 9 = \frac 12\] first step is to change it into exponential form \[x^{1/2} = 9\] do you agree with this form

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

good. so we need x. you have x^1/2 what do you think you should do to make that x^1/2 into just x^1?

OpenStudy (anonymous):

could you log both sides, and bring down the 1/2

OpenStudy (lgbasallote):

hmm not exactly

OpenStudy (lgbasallote):

here's a hint \[\LARGE (a^3)^{1/3} = a^{3/3} = a^1 = a\] \[\LARGE (a^2)^{1/2} = a^{2/2} = a^1 = a\] \[\LARGE (a^{1/3})^3 = a^{3/3} = a^1 = a\]

OpenStudy (anonymous):

oh, do you square both sides?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

so x = 81?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

perfect thanks!